# DRAT Proofs of Unsatisfiability for SAT Modulo Monotonic Theories

### 19 Jan 2024 v1

@@ 6 Experimental Evaluation
To evaluate our proposed method, we implemented it as shown earlier in Fig. 1 (Sec. 3). We call our implementation MonoProof [^4] All source files and benchmarks are available in https://github.com/NickF0211/MonoProof..
@@ 6.1 Benchmarks
We wish to evaluate scalability on real, industrial problems arising in practice. MonoProof has successfully generated and verified industrial UNSAT proofs for a set of hard, unsatisfiable Tiros [2, 8] queries collected in production use at AWS over a multi-week period. However, these instances are proprietary and cannot be published, making them irreproducible by others. Instead, we evaluate on two sets of benchmarks that we can publicly release:
@@ 6.2 Results
2. MonoProof’s time to prepare and check a proof of unsatisfiability is markedly faster than standard bit-blasting or lemma-specific bit-blasting. Fig. 3 summarizes our results (full table in Appendix 0.D). The left graph shows solving times (with proof logging). Since the proof-logging overhead is so low for both bit-blasting (Kissat generating DRAT) and MonoProof, these results are consistent with prior work showing the superiority of the SMMT approach for solving [9]. Note that bit-blasting (BB) solved all 24 network reachability instances, but failed to solve any of the 24 escape routing instances in the 1hr timeout. Lemma-specific bit-blasting (LSBB) and MonoProof share the same solving and proof-logging steps. The right graph shows proof-checking times (including BackwardCheck and proof-specific Horn upper-bound construction for MonoProof). Here, BB could proof-check only 11/24 reachability instances that it had solved. Restricting to only the 11 instances that BB proof-checked, MonoProof was at least 3.7×3.7$\times$3.7 × and geometric mean (GM) 10.2×10.2$\times$10.2 × faster. LSBB timed out on all 48 instances. Summarizing, MonoProof solved and proved all 48 instances, whereas BB managed only 11 instances, and LSBB failed to prove any.
@@ 7 Conclusion
Appendix 0.A Appendix: Supplementary Proofs
0.A.1 Proof of Theorem 4.2 (Monotonic Transformation)
@@ Appendix 0.A Appendix: Supplementary Proofs
Let $\Sigma_{\mathbf{p}}^{+}$ be the result of applying monotonic transformation MONOT on a propositional definition $\Sigma_{\mathbf{p}}$ of $\mathbf{p}$ where $\mathbf{p}$ is the positive predicate atom for monotonic predicate $p$ over inputs $A$. We will now prove that $\Sigma_{\mathbf{p}}^{+}$ is a propositional definition for $\mathbf{p}$. First, for any assignment $\mathcal{M}$ over $A$, $\Sigma_{\mathbf{p}}^{+}|_{M}$ is satisfiable because $\Sigma_{\mathbf{p}}|_{M}$ is satisfiable. We can obtain a satisfying solution to $\Sigma_{\mathbf{p}}^{+}|_{M}$ by extending a solution to $\Sigma_{\mathbf{p}}|_{M}$ by assigning $\mathbf{a^{\prime}}$ and $\mathbf{p}$ to the same truth values as $\mathbf{a}$ and $\mathbf{p}$, respectively. Second, we prove, $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$ if and only if $p(\mathcal{M}(A))=\top$. For the forward direction, if $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$, then it must be the case $\Sigma_{\mathbf{p}}\models(\mathcal{M}\Rightarrow\mathbf{p})$ since there is $\Sigma_{\mathbf{p}}^{+}$ is an upper-bound of $\Sigma_{\mathbf{p}}$. Therefore, we can conclude $p(\mathcal{M}(A))=\top$. For the backward direction, if $p(\mathcal{M}(A))=\top$, and then $\Sigma_{\mathbf{p}}\models(\mathcal{M}\Rightarrow\mathbf{p})$. Since $p$ is monotonic, we have $\Sigma_{\mathbf{p}}\models(\mathcal{M}_{A}\Rightarrow\mathbf{p})$ where $\mathcal{M}_{A}$ is the set of positive input atoms in $\mathcal{M}$. Therefore, $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}_{A}^{\prime}\Rightarrow\mathbf{p^{\prime}})$ where $\mathcal{M}_{A}^{\prime}=\{\mathbf{a^{\prime}}\mid\mathbf{a}\in\mathcal{M}_{A}\}$. Since $\Sigma_{\mathbf{p}}^{+}$ contains the Horn clauses: $\mathbf{a}\Rightarrow{\mathbf{a^{\prime}}}$ for every $a\in A$ and $\mathbf{p^{\prime}}\Rightarrow\mathbf{p}$, we have $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}_{A}\Rightarrow\mathbf{p})$. Since $\mathcal{M}\supseteq\mathcal{M}_{A}$, we conclude $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$.
@@ 0.A.2 Proof of Lemma 1 (Duality)
First, we prove: $\Sigma_{\mathbf{p}}\models(M\Rightarrow\mathbf{p})$ if and only if $\Sigma_{\bar{\mathbf{p}}}\wedge M\wedge\mathbf{p}$. By Def. 2, if $p(M)$ evaluates to $\top$, then $\Sigma_{\mathbf{p}}\models(M\Rightarrow\mathbf{p})$. Therefore, $\Sigma_{\bar{\mathbf{p}}}\not\models(M\Rightarrow\bar{\mathbf{p})}$, and hence $\Sigma_{\bar{\mathbf{p}}}\wedge M\wedge\mathbf{p}$ is satisfiable.
The second statement in Lemma 1, $\Sigma_{\bar{\mathbf{p}}}\models(M\Rightarrow\bar{\mathbf{p})}$ if and only if $\Sigma_{\mathbf{p}}\wedge M\wedge\bar{\mathbf{p}}$, can be proved analogously. ∎
@@ 0.A.3 Proof of Lemma 2 (Instantiation-Based Upper-Bound)
To prove $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\Rightarrow\overline{\mathbf{p}}$ is an upper-bound of $\Sigma_{\bar{\mathbf{p}}}$, it is sufficient to show that for any theory lemma $M_{A}\Rightarrow\overline{\mathbf{p}}$ that is logically implied by $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\overline{\mathbf{p}}$, the lemma is also logically implied by $\Sigma_{\overline{\mathbf{p}}}$. If $M_{A}\Rightarrow\mathbf{\overline{p}}$ is logically implied by $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\mathbf{\overline{p}}$, then the left-hand side of $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\mathbf{\overline{p}}$ (i.e., $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}$) must be satisfiable under the assignment $\neg(M_{A}\Rightarrow\overline{\mathbf{p}})$. Therefore, $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\wedge\neg(M_{A}\Rightarrow\bar{\mathbf{p}})$ must be satisfiable, and thus $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\wedge M_{A}\wedge\mathbf{p}$ is satisfiable. Since the variables in $M^{\prime}$ are disjoint from the variables in $M_{A}$ and $var(\mathbf{p})$, the satisfibility of $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\bar{p}}}\wedge M_{A}\wedge\mathbf{p}$ implies that $\Sigma_{\mathbf{p}}|_{\mathbf{\bar{p}}}\wedge M^{\prime}\wedge M_{A}\wedge\mathbf{p}$ is satisfiable. Moreover, since $\bar{\mathbf{p}}$ is not a variable in $\Sigma_{\mathbf{p}}|_{\mathbf{\bar{p}}}$, the satisfibility of $\Sigma_{\mathbf{p}}|_{\mathbf{\bar{p}}}\wedge M^{\prime}\wedge M_{A}\wedge\mathbf{p}$ implies the satisfiability of $\Sigma_{\mathbf{p}}|_{\mathbf{\bar{p}}}\wedge M^{\prime}\wedge M_{A}$, and it further implies the satisfiability of $\Sigma_{\mathbf{p}}\wedge\mathbf{\overline{p}}\wedge M^{\prime}\wedge M_{A}$ since $\mathbf{p}$ does not appear in $M^{\prime}$ nor $M$. Therefore, by dropping the assignment $M^{\prime}$ on the satisfiable formula $\Sigma_{\mathbf{p}}\wedge\mathbf{\overline{p}}\wedge M^{\prime}\wedge M_{A}$, we obtain another satisfiable formula $\Sigma_{\mathbf{p}}\wedge\mathbf{\bar{p}}\wedge M_{A}$. By the second condition of Duality (Lemma. 1), $M_{A}\Rightarrow\bar{\mathbf{p}}$ is logically implied by $\Sigma_{\bar{\mathbf{p}}}$. ∎
@@ 0.A.4 Proof of Lemma 3 (Instantiation-Based Horn Upper-Bound)
By lemma 2, $(\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\bar{\mathbf{p}}$ is an upper-bound of $\Sigma_{\bar{\mathbf{p}}}^{+}$. Moreover, since $\Sigma_{\mathbf{p}}^{+}$ is a monotonic definition, $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}}$ contains only negative literals of $A$, the encoding of $(\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\bar{\mathbf{p}}$ contains two type of clauses: (1) $\mathbf{a}\vee\bar{c_{i}}$ and (2) $c_{1}\vee c_{2}\ldots c_{n}\vee\bar{\mathbf{p}}$ where $c_{i}$ is a fresh literal introduced for every clause of $\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}}$. Therefore, the formula $(\Sigma_{\mathbf{p}}^{+}|_{\bar{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\bar{\mathbf{p}}$ is dual-Horn.
0.A.5 Proof of Theorem 4.3 (Instantiation-Based Lemma-Specific Definition)

### 04 Apr 2024 v2

@@ 6 Experimental Evaluation
To evaluate our proposed method, we implemented it as shown earlier in Fig. 1 (Sec. 3). We call our implementation MonoProof (available at https://github.com/NickF0211/MonoProof).
@@ 6.1 Benchmarks
We wish to evaluate scalability on real, industrial problems arising in practice. MonoProof has successfully generated and verified industrial UNSAT proofs for a set of hard, unsatisfiable Tiros [2, 8] queries collected in production use at AWS over a multi-week period. However, these instances are proprietary and cannot be published, making them irreproducible by others. Instead, we evaluate on two sets of benchmarks that we can publicly release (also at https://github.com/NickF0211/MonoProof):
@@ 6.2 Results
[…]
@@ 7 Conclusion
Acknowledgments
Nick Feng was supported in part by an Amazon Research Award. Alan Hu was supported in part by a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada (NSERC). The authors would like to thank Marijn Heule for insightful feedback and mentorship during Nick’s internship at Amazon Web Services in 2021. We also thank Dan Dacosta, Nadia Labai, and Nate Launchbury for reviewing earlier drafts of this work.
Appendix 0.A Supplementary Proofs
0.A.1 Proof of Theorem 4.2 (Correctness of Monotonic Transformation)
To improve readability, we repeat the key definitions/theorems here:
Definition 2 (Propositional Definition).
Definition 5 (Monotonic Definition).
Definition 6 (Monotonic Transformation).
Theorem 4.2 (Correctness of Monotonic Transformation).
The intuition behind the proof is that the requirements to be a monotonic definition are specified in terms of only the input atoms $A$ and the predicate atom $\mathbf{p}$. Because the predicate $p$ is monotonic, and our propositional definitions need only be one-sided, it’s possible to “hide”, via renaming variables, any clauses that violate the requirements to be a monotonic definition.
@@ Appendix 0.A Supplementary Proofs
It’s obvious that $\textsc{MonoT}(\mathbf{p},\Sigma_{\mathbf{p}})$ results in $n+|A|+1$ clauses: the original $n$ clauses in $\Sigma_{\mathbf{p}}$ (with variables renamed as specified), plus the $|A|+1$ new implications specified in the construction in Def. 6. The main goal in the proof is to establish that $\Sigma_{\mathbf{p}}^{+}$ is a monotonic definition.
Referring to Def. 5, conditions (2) and (3) are trivially satisfied: $\Sigma_{\mathbf{p}}^{+}$ does not contain $\mathbf{\overline{p}}$ because of the variable renaming, and $\mathbf{p}$ appears only in the Horn clause $\mathbf{p^{\prime}}\Rightarrow\mathbf{p}$. So what remains is to establish that $\Sigma_{\mathbf{p}}^{+}$ is a propositional definition for $\mathbf{p}$.
First, for any assignment $\mathcal{M}$ over $A$, $\Sigma_{\mathbf{p}}^{+}|_{M}$ is satisfiable because $\Sigma_{\mathbf{p}}|_{M}$ is satisfiable. We can obtain a satisfying solution to $\Sigma_{\mathbf{p}}^{+}|_{M}$ by extending a solution to $\Sigma_{\mathbf{p}}|_{M}$ by assigning $\mathbf{a^{\prime}}$ and $\mathbf{p^{\prime}}$ to the same truth values as $\mathbf{a}$ and $\mathbf{p}$, respectively.
Second, for any assignment $\mathcal{M}$ over $A$, we prove that $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$ if and only if $p(\mathcal{M})=\top$:
  • For the forward direction, if $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$, then we will prove that $p(\mathcal{M})=\top$ by contradiction. Assume that there is a $\mathcal{M}$ such that $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$ and $p(\mathcal{M})=\bot$. Since $\Sigma_{\mathbf{p}}$ is a propositional definition of $\mathbf{p}$, by Def. 2, we know from $p(\mathcal{M})=\bot$ that $\Sigma_{\mathbf{p}}\not\models(\mathcal{M}\Rightarrow\mathbf{p})$. Therefore, there exists a satisfying solution $M^{\prime}$ to the formula $\Sigma_{\mathbf{p}}\wedge\mathcal{M}\wedge\mathbf{p}$. We can extend $M^{\prime}$ to obtain a satisfying assignment $M^{\prime+}$ to the formula $\Sigma_{\mathbf{p}}^{+}\wedge\mathcal{M}\wedge\mathbf{p}$ by assigning $\mathbf{a^{\prime}}$ and $\mathbf{p^{\prime}}$ to the same truth values as $\mathbf{a}$ and $\mathbf{p}$. Therefore, $\Sigma_{\mathbf{p}}^{+}\not\models(\mathcal{M}\Rightarrow\mathbf{p})$, which is a contradiction.
  • For the backward direction, if $p(\mathcal{M})=\top$, then $\Sigma_{\mathbf{p}}\models(\mathcal{M}\Rightarrow\mathbf{p})$. Since $p$ is monotonic, we have $\Sigma_{\mathbf{p}}\models(\mathcal{M}_{A}\Rightarrow\mathbf{p})$ where $\mathcal{M}_{A}$ is the set of positive input atoms in $\mathcal{M}$. Therefore, $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}_{A}^{\prime}\Rightarrow\mathbf{p^{\prime}})$ where $\mathcal{M}_{A}^{\prime}=\{\mathbf{a^{\prime}}\mid\mathbf{a}\in\mathcal{M}_{A}\}$. Since $\Sigma_{\mathbf{p}}^{+}$ contains the Horn clauses: $\mathbf{a}\Rightarrow{\mathbf{a^{\prime}}}$ for every $a\in A$ and $\mathbf{p^{\prime}}\Rightarrow\mathbf{p}$, we have $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}_{A}\Rightarrow\mathbf{p})$. Since $\mathcal{M}\supseteq\mathcal{M}_{A}$, we conclude $\Sigma_{\mathbf{p}}^{+}\models(\mathcal{M}\Rightarrow\mathbf{p})$.
@@ 0.A.2 Proof of Lemma 1 (Duality)
Lemma 1 (Duality).
Let $p$ be a monotonic predicate over Boolean arguments $A$. Suppose $\Sigma_{\mathbf{p}}$ and $\Sigma_{\overline{\mathbf{p}}}$ are positive and negative propositional definitions, respectively. For every assignment $\mathcal{M}$ to the variables in $A$:
  1. 1.
    $\Sigma_{\mathbf{p}}\models(\mathcal{M}\Rightarrow\mathbf{p})$ if and only if $\Sigma_{\overline{\mathbf{p}}}\wedge\mathcal{M}\wedge\mathbf{p}$ is satisfiable.
  2. 2.
    $\Sigma_{\overline{\mathbf{p}}}\models(\mathcal{M}\Rightarrow\overline{\mathbf{p}})$ if and only if $\Sigma_{\mathbf{p}}\wedge\mathcal{M}\wedge\overline{\mathbf{p}}$ is satisfiable.
@@
To prove $\Sigma_{\mathbf{p}}\models(M\Rightarrow\mathbf{p})$ if and only if $\Sigma_{\overline{\mathbf{p}}}\wedge M\wedge\mathbf{p}$ is satisfiable:
$\Sigma_{\mathbf{p}}\models(M\Rightarrow\mathbf{p})$ iff $p(M)=\top$ By Def. 2 since $\Sigma_{\mathbf{p}}$ is a
propositional definition.
iff $p(M)\not=\bot$
iff $\Sigma_{\overline{\mathbf{p}}}\not\models(M\Rightarrow\overline{\mathbf{p}})$ By Def. 2.
iff $\Sigma_{\overline{\mathbf{p}}}\wedge\overline{(M\Rightarrow\overline{\mathbf{p}})}$ is satisfiable By definition of $\models$.
iff $\Sigma_{\overline{\mathbf{p}}}\wedge M\wedge\mathbf{p}$ is satisfiable
The second statement in Lemma 1, $\Sigma_{\overline{\mathbf{p}}}\models(M\Rightarrow\overline{\mathbf{p})}$ if and only if $\Sigma_{\mathbf{p}}\wedge M\wedge\overline{\mathbf{p}}$, can be proved analogously. ∎
Lemma 1 illuminates a connection between any positive and negative propositional definitions $\Sigma_{\mathbf{p}}$ and $\Sigma_{\overline{\mathbf{p}}}$, even those these definitions are one-sided.
@@ 0.A.3 Proof of Lemma 2 (Instantiation-Based Upper-Bound)
Lemma 2 (Instantiation-Based Upper-Bound).
Let a predicate $p$ over input $A$ and a positive definition $\Sigma_{\mathbf{p}}$ be given. For any partial assignment $M^{\prime}$ over $var(\Sigma_{\mathbf{p}})\setminus(var(p)\cup A)$,   $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\overline{\mathbf{p}}$ is an over-approximation of $\Sigma_{\overline{\mathbf{p}}}$.
@@
To prove that $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\overline{\mathbf{p}}$ is an over-approximation of any propositional definition $\Sigma_{\overline{\mathbf{p}}}$, we will show that for any theory lemma $M_{A}\Rightarrow\overline{\mathbf{p}}$ that is logically implied by $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\overline{\mathbf{p}}$, the lemma is also logically implied by $\Sigma_{\overline{\mathbf{p}}}$.
If the lemma $M_{A}\Rightarrow\mathbf{\overline{p}}$ is logically implied by $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\mathbf{\overline{p}}$, then
$\left(\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\mathbf{\overline{p}}\right)\wedge\neg\left(M_{A}\Rightarrow\mathbf{\overline{p}}\right)$ (1)
must be false (i.e., UNSAT), by the definition of implication. Note, however, that $\neg\left(M_{A}\Rightarrow\mathbf{\overline{p}}\right)$ is just the conjunction of literals (unit-clauses) $M_{A}\wedge\mathbf{p}$, which we can interpret as an assignment.
Now, consider applying that assignment $M_{A}\wedge\mathbf{p}$ to $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}$, namely:
$\left(\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\right)|_{M_{A}\cup\mathbf{p}}.$ (2)
We can see that statement 2 must be valid, because if it had any falsifying assignment, that assignment would make $\left(\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\mathbf{\overline{p}}\right)$ true, so we could extend that assignment with $M_{A}\cup\mathbf{p}$ and satisfy statement 1, which we know must be UNSAT. Therefore, since we have established that $\left(\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\right)|_{M_{A}\cup\mathbf{p}}$ is valid, we know that
$\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\wedge M_{A}\wedge\mathbf{p}$ (3)
must be satisfiable (using the assignment $M_{A}\cup\mathbf{p}$ and any assignment to the remaining variables).
From the satisfiability of statement 3, we derive that $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\wedge M_{A}$ is satisfiable (since omitting a conjunct can only increase satisfiability). Furthermore, since the variables in $M^{\prime}$ are disjoint from the variables in $M_{A}$ and $var(\mathbf{p})$, there can be no conflicts between these partial assignments, so the satisfiability of $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\wedge M_{A}$ implies the satisfiability of $\Sigma_{\mathbf{p}}\wedge M^{\prime}\wedge\mathbf{\overline{p}}\wedge M_{A}$. Weakening again by omitting conjuncts, we obtain that $\Sigma_{\mathbf{p}}\wedge\mathbf{\bar{p}}\wedge M_{A}$ is satisfiable. By the second condition of Duality (Lemma 1), $M_{A}\Rightarrow\overline{\mathbf{p}}$ is logically implied by $\Sigma_{\bar{\mathbf{p}}}$. ∎
Lemma 2 establishes that arbitrary partial assignments to a positive propositional definition can be used to create a safe upper-bound of a negative propositional definition, and vice-versa. But the lemma does not establish that these upper-bounds are particularly tight. In Theorem 4.3, we show that it is always possible to carefully choose an assignment to create an upper-bound of a propositional definition sufficient to prove any specific theory lemma.
@@ 0.A.4 Proof of Lemma 3 (Instantiation-Based Horn Upper-Bound)
To improve readability, we repeat a key definition and lemma here:
Definition 5 (Monotonic Definition).
Lemma 2 (Instantiation-Based Upper-Bound).
Let a predicate $p$ over input $A$ and a positive definition $\Sigma_{\mathbf{p}}$ be given. For any partial assignment $M^{\prime}$ over $var(\Sigma_{\mathbf{p}})\setminus(var(p)\cup A)$,   $\Sigma_{\mathbf{p}}|_{M^{\prime}\cup\mathbf{\overline{p}}}\Rightarrow\overline{\mathbf{p}}$ is an over-approximation of $\Sigma_{\overline{\mathbf{p}}}$.
It’s also important to reiterate the text of a footnote originally attached to Lemma 2:
Note that $\Sigma_{\mathbf{p}}|_{M^{\prime}}$ is encoded in CNF, so to compactly (i.e., linear-size) encode $\Sigma_{\mathbf{p}}|_{M^{\prime}}\Rightarrow\overline{\mathbf{p}}$ in CNF, we introduce a new literal $l_{i}$ for each clause $C_{i}\in\Sigma_{\mathbf{p}}|_{M^{\prime}}$, create clauses $\overline{c_{ij}}\vee\overline{l_{i}}$ for each literal $c_{ij}\in C_{i}$, and add clause $l_{1}\vee l_{2}\vee\ldots\vee l_{n}\vee\overline{\mathbf{p}}$.
which describes how we encode $\Sigma_{\mathbf{p}}|_{M^{\prime}}\Rightarrow\overline{\mathbf{p}}$ into CNF. We use Lemma 2 with this encoding in Lemma 3, where we need to establish that these new clauses are all (dual) Horn. (The original clauses in $\Sigma_{\mathbf{p}}|_{M^{\prime}}$ are not included in the newly encoded CNF.)
Lemma 3 (Instantiation-Based Horn Upper-Bound).
Given a monotonic predicate $p$ over input $A$ and a positive monotonic definition $\Sigma_{\mathbf{p}}^{+}$, let $X$ represent the set of auxiliary variables: $var(\Sigma_{\mathbf{p}}^{+})\setminus(A\cup var(p))$. For any complete satisfying assignment $M_{X\cup A}$ to $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}}$, the formula $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\overline{\mathbf{p}}$ serves as a Horn upper-bound for any propositional definition of $\overline{\mathbf{p}}$, where $\mathcal{M}_{X}$ is a partial assignment derived from $M_{X\cup A}$ for the auxiliary variables $X$.
@@
By lemma 2, $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\overline{\mathbf{p}}$ is an upper-bound of $\Sigma_{\overline{\mathbf{p}}}^{+}$. All that remains to be established is that this is a Horn upper-bound.
Since $M_{X}$ is a complete assignment to the auxiliary variables, the clauses in $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})$ can contain only the variables in $A$. Furthermore, by Def. 5, all literals for variables in $A$ are negative.
The CNF that we build for $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\overline{\mathbf{p}}$, consists entirely of two types of clauses: (1) clauses of the form $(\overline{c_{ij}}\vee\overline{l_{i}})$ for each literal $c_{ij}$ in each original clause $C_{i}\in(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})$, and (2) a single clause $(l_{1}\vee l_{2}\vee\ldots\vee l_{n}\vee\overline{\mathbf{p}})$. The preceding paragraph established that all the literals $c_{ij}$ are negative literals, so the literals $\overline{c_{ij}}$ are positive. Therefore, all clauses of type (1) are dual Horn. (They are also Horn, but that’s irrelevant for this proof.) The clause of type (2) is obviously dual Horn. Therefore, the CNF we construct for the formula $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\overline{\mathbf{p}}$ is dual Horn. ∎
0.A.5 Proof of Theorem 4.3 (Lemma-Specific Horn Upper-Bound)
To improve readability, we repeat key definitions and lemmas here:
Definition 2 (Propositional Definition).
Definition 5 (Monotonic Definition).
Lemma 3 (Instantiation-Based Horn Upper-Bound).
Given a monotonic predicate $p$ over input $A$ and a positive monotonic definition $\Sigma_{\mathbf{p}}^{+}$, let $X$ represent the set of auxiliary variables: $var(\Sigma_{\mathbf{p}}^{+})\setminus(A\cup var(p))$. For any complete satisfying assignment $M_{X\cup A}$ to $\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}}$, the formula $(\Sigma_{\mathbf{p}}^{+}|_{\overline{\mathbf{p}}\cup\mathcal{M}_{X}})\Rightarrow\overline{\mathbf{p}}$ serves as a Horn upper-bound for any propositional definition of $\overline{\mathbf{p}}$, where $\mathcal{M}_{X}$ is a partial assignment derived from $M_{X\cup A}$ for the auxiliary variables $X$.
Theorem 4.3 (Lemma-Specific Horn Upper-Bound).